# BLAZE: Cross-Language and Cross-Project Bug Localization via Dynamic Chunking and Hard Example Learning

### v2

@@ VI Threats to Validity
Internal validity. One threat to internal validity is the hyperparameters we used to train the models. To mitigate this, we use the same parameters used by several prior studies that use large language models [36, 19, 62, 13].
External validity. A threat to the external validity of our study lies in the generalizability of the results. We assess cross-project capabilities using separate projects for training and testing. Although the choice of test projects may affect performance, this approach aligns with previous studies that evaluate zero-shot and cross-project capabilities [58, 32]. Additionally, our evaluation of 31,745 real-world bugs ensures that our findings hold, even with different project selections, due to the large scale of our dataset. To address potential selection bias in the BeetleBox dataset, we collected project repositories based on star ratings and recent updates, ensuring that the repositories are actively contributed to and popular within the open-source community. This approach aims to include high-quality, widely used projects across the selected top five programming languages. Additionally, due to the absence of a replication package, we replicated TRANP-CNN and CooBA based on their descriptions in the original study, which may lead to slight performance deviations. Nevertheless, after experimenting with various hyperparameters, we selected a set that achieves comparable performance to that reported in the original study.
@@ VII Related Works
A plethora of work proposed lexical retriever-based approaches to localizing bugs. Such approaches use BM25 and TF-IDF techniques to measure the similarity between a bug report and source code files [5, 6, 64, 65, 66, 67]. For example, BugLocator [5] used TF-IDF to measure the similarity between the bug report and source code files using a revised vector space model. Sisman et al. [68] and Amalgam [6] used version history along with the VSM model to localize bugs. Shi et al. [69] found that BM25 and its variants are better than TF-IDF, VSM, and rVSM in the bug localization task. BoostNSift [65] used BM25 for method-level bug localization. Our approach, BLAZE, differs from prior work since it leverages the advantages of both the BM25-based retrieval method and deep learning-based technique to perform cross-project bug localization.
Other work proposed deep learning-based approaches to localizing bugs [63, 52, 5, 70, 11, 71]. For example, FLIM [63] used CodeBERT embedding for calculating the similarity between bug reports and source code files at the function level. DNNLoc [72] combined a deep neural network with the VSM model to increase effectiveness across different types of metrics such as textual similarity, class name similarity, bug-fixing recency, and frequency. NP-CNN [52] employed a Convolutional Neural Network (CNN) to leverage both lexical and semantical properties. BL-GAN [12] proposed a generative adversarial network-based solution for bug localization. Several studies [51, 9, 8, 10] incorporated program structure information in the model to generate more useful program representation. For example, Liang et al. [10] and Li et al. [8] used abstract syntax trees, data flow graphs, and program flow graphs to generate a better representation of source code. While BLAZE employs an embedding model similar to the techniques discussed previously, it differs by its ability to be applied to new, unseen projects without requiring additional training.
@@ VIII Conclusion
In this paper, we introduce BLAZE, a bug localization tool that operates without project-specific training, facilitating cross-project and cross-language bug localization. BLAZE employs a dynamic chunking and hard example learning methodology to model lengthy source code files. To support cross-project and cross-language bug localizers, we curate BeetleBox—one of the largest (if not the largest) cross-project and cross-language bug localization datasets. We evaluate BLAZE with 31,745 bug reports from 34 real-world projects, showing up to a 100% improvement in the Top 1 metric against cross-project localizers and up to a 60% improvement over modern large language model approaches. Our results suggest that the use of dynamic chunking and hard example learning is key in enhancing the performance of bug localization tools.

### v3

@@ VI Threats to Validity
Internal validity. One threat to internal validity is the hyperparameters we used to train the models. To mitigate this, we use the same parameters used by several prior studies that use large language models [36, 19, 63, 13]. Although we applied conservative repository-level filtering, we recognize that some residual risk of leakage could persist due to indirect or partial content overlaps. However, given the lack of transparency in LLM training corpora, our approach represents a reasonable trade-off between feasibility and rigor. Furthermore, during the filtering process, we exclude bug reports that explicitly mention modified files. However, subtle or indirect references may still go undetected. This presents a residual threat to internal validity, as a small number of answer-revealing reports may persist despite our conservative heuristics.
External validity. A threat to the external validity of our study lies in the generalizability of the results. We assess cross-project capabilities using separate projects for training and testing. Although the choice of test projects may affect performance, this approach aligns with previous studies that evaluate zero-shot and cross-project capabilities [58, 32]. Additionally, our evaluation of 30,625 real-world bugs ensures that our findings hold, even with different project selections, due to the large scale of our dataset. To address potential selection bias in the BeetleBox dataset, we collected project repositories based on star ratings and recent updates, ensuring that the repositories are actively contributed to and popular within the open-source community. This approach aims to include high-quality, widely used projects across the selected top five programming languages. Additionally, due to the absence of a replication package, we replicated TRANP-CNN and CooBA based on their descriptions in the original study, which may lead to slight performance deviations. Nevertheless, after experimenting with various hyperparameters, we selected a set that achieves comparable performance to that reported in the original study.
@@ VII Related Works
A plethora of work proposed lexical retriever-based approaches to localizing bugs. Such approaches use BM25 and TF-IDF techniques to measure the similarity between a bug report and source code files [5, 6, 65, 66, 67, 68]. For example, BugLocator [5] used TF-IDF to measure the similarity between the bug report and source code files using a revised vector space model. Sisman et al. [69] and Amalgam [6] used version history along with the VSM model to localize bugs. Shi et al. [70] found that BM25 and its variants are better than TF-IDF, VSM, and rVSM in the bug localization task. BoostNSift [66] used BM25 for method-level bug localization. Lee et al. [55] proposed the Bench4BL dataset and evaluated six bug-localization tools. Our study differs from Lee et al.’s, mainly in the reliability of ground truth mappings. While Lee et al. used the Bench4BL dataset—later shown by Kim et al. [56] to have up to 44% incorrect mappings—we applied Kim et al.’s method to ensure accuracy, resulting in only 0.06% incorrect mappings. This improvement is likely due to our conservative issue selection, including only issues confirmed to be merged, unlike in Lee et al.’s approach. Our approach, BLAZE, differs from prior work since it leverages the advantages of both the BM25-based retrieval method and deep learning-based technique to perform cross-project bug localization.
Other work proposed deep learning-based approaches to localizing bugs [64, 52, 5, 71, 11, 72, 73]. For example, FLIM [64] used CodeBERT embedding for calculating the similarity between bug reports and source code files at the function level. DNNLoc [74] combined a deep neural network with the VSM model to increase effectiveness across different types of metrics such as textual similarity, class name similarity, bug-fixing recency, and frequency. NP-CNN [52] employed a Convolutional Neural Network (CNN) to leverage both lexical and semantical properties. BL-GAN [12] proposed a generative adversarial network-based solution for bug localization. Several studies [51, 9, 8, 10] incorporated program structure information in the model to generate more useful program representation. For example, Liang et al. [10] and Li et al. [8] used abstract syntax trees, data flow graphs, and program flow graphs to generate a better representation of source code. While BLAZE employs an embedding model similar to the techniques discussed previously, it differs by its ability to be applied to new, unseen projects without requiring additional training.
@@ VIII Conclusion
In this paper, we introduce BLAZE, a bug localization tool that operates without project-specific training, facilitating cross-project and cross-language bug localization. BLAZE employs a dynamic chunking and hard example learning methodology to model lengthy source code files. To support cross-project and cross-language bug localizers, we curate BeetleBox—one of the largest (if not the largest) cross-project and cross-language bug localization datasets. We evaluate BLAZE with 30,625 bug reports from 34 real-world projects, showing up to a 100% improvement in the Top 1 metric against cross-project localizers and up to a 60% improvement over modern large language model approaches. Our results suggest that the use of dynamic chunking and hard example learning is key in enhancing the performance of bug localization tools.